A shirt was originally priced at $80. The store is having a 15% off sale. How much will you pay for the shirt after the discount (not including tax)?
Answer by ikleyn(52781)

.

    80 - 0.15*80 = (1-0.15)*80 = 0.85*80 = 68 dollars.    ANSWER

Solved.